A student ran the following reaction in the laboratory at 294 K:

2CH2Cl2(g) ____> CH4(g) + CCl4(g)

When she introduced 7.04×10-2 moles of CH2Cl2(g) into a 1.00 Liter container, she found the equilibrium concentration of CCl4(g) to be 3.25×10-2 M.

Calculate the equilibrium constant, Kc, she obtained for this reaction.
Answer #1

Ans :- Equilibrium constant i.e. kc = 3.62 x 10-1

Explanation :-

Given,

Initial number of moles of CH2Cl2 = 7.04 x 10-2 mol

Volume of the container = 1.00 L

As,

Concentration of Substance = Number of moles of solute / Volume of the solution in L

So,

Initial concentration of CH2Cl2 = Number of moles of CH2Cl2 / Volume of the solution in L

= 7.04 x 10-2 mol / 1 L

= 7.04 x 10-2 M

ICE table is :

...........................2 CH2Cl2 (g) <-----------------------------> CH4 (g)...............+...................CCl4 (g)

Initial (I)................7.04 x 10-2 M........................................0.0 M.......................................0.0 M

Change (C).............-2y ......................................................+y............................................+y

Equilibrium (E)....(7.04 x 10-2 M - 2y) M ................................y M.........................................y M

Here,

y = Amount dissociated per mole

Given,

Equilibrium concentration of CCl4 (g) = 3.25 x 10-2 M = y

So,

y = 3.25 x 10-2  

Also,

Equilibrium concentration of CH2Cl2 (g) = (7.04 x 10-2 M - 2y) M = (7.04 x 10-2 - 2x3.25 x 10-2 ) M = 5.4 x 10-3 M

Equilibrium concentration of CH4 (g) = Equilibrium concentration of CCl4 (g) = y = 3.25 x 10-2 M

Expression of Equilibrium constant i.e. Kc(which is equal to the product of the molar concentration of products divided by product of the molar concentration of reactants raise to power their stoichiometric coefficient when reaction is at equilibrium stage) is.

kc = [CH4].[CCl4] / [CH2Cl2]2

kc = (3.25 x 10-2 M)2 / (5.4 x 10-3 M)2

kc = 0.362 = 3.62 x 10-1

Hence, Equilibrium constant = kc = 3.62 x 10-1